A woman invested 4 times as much at 4% as she did at 7%. The total amount of interest she earns in 1 year from both accounts is $2,070. How much did she invest at each rate?

Answer by mananth(16949)

7.00% per annum ------------- x
4.00% per annum ------------ y 4 x

Interest----- $2,070.00

7.00% x + 4.00% * 4 x = $2,070.00
0.07 x + 0.04 * 4 x = $2,070.00
0.07 x + 0.16 x = $2,070.00
0.23 x = $2,070.00
/ 0.23
x= $9,000.00 @ 7%
4 x = $36,000.00 @4%

CHECK
$630.00 + $1,440.00 = $2,070.00
m.ananth@hotmail.ca